# Double Robust Variance Estimation

### v1

@@ 2.3.2 Empirical sandwich variance estimator
where $\psi_{\gamma}$ is the vector of score functions for the scaled outcome model. The estimating functions $\psi_{t1}$ and $\psi_{t0}$ are for the two targeting models, while $\psi_{\mu_{TMLE1}}$ and $\psi_{\mu_{TMLE0}}$ are for the causal means $\mu_{TMLE}^{1}$ and $\mu_{TMLE}^{0}$, respectively. Note $c_{x}(O_{i},\gamma,\eta_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\gamma)\}+\eta_{x}](b-a)+a$ are the pseudo-outcomes following the targeting step, as outlined in Section 2.2.3. Finally, $\psi_{TMLE}$ is a delta method transformation of the causal means for estimation of the $ACE$ using $\widehat{DR}_{TMLE}$. Unbiasedness of $\psi_{3}(O_{i};\theta)$ when at least one nuisance model is correctly specified is demonstrated in the Appendix.
[…]
3 Example: Improving Pregnancy Outcomes with Progesterone
Prior research has shown links between maternal anemia and low birth weight (Azizah et al.,, 2022; Figueiredo et al.,, 2018). The relationship between anemia and birth weight is particularly important to understand for women living with HIV, as they experience anemia more frequently than women without HIV (Levine et al.,, 2001) and are more prone to adverse birth outcomes (Tukei et al.,, 2021). The effect of maternal anemia on birth weight among women living with HIV was estimated with data from the IPOP trial.
The IPOP trial was a randomized, double-blind, placebo-controlled trial conducted in Lusaka, Zambia. Participants were pregnant women aged 18 or older living with HIV on (or intending to initiate) antiretroviral therapy (ART) who had viable singleton pregnancies at less than 24 weeks of gestation at baseline. The 800 enrolled participants were randomized with equal probability to either a placebo or 17 alpha-hydroxyprogesterone caproate (17P) treatment to evaluate the efficacy of 17P for prevention of preterm birth. Additional details about the IPOP trial have been published previously (Price et al.,, 2021). While IPOP was designed to examine the effect of 17P on preterm birth or stillbirth, covariates collected during the trial allow for examination of other causal questions, assuming that the identification assumptions hold.
The three doubly robust estimators described in Section 2.2 were applied to IPOP data for estimation of the $ACE$ of maternal anemia on birth weight. The exposure, maternal anemia, was defined as having a baseline hemoglobin level below 10.5 g/dL. The IPOP data were limited to the 782 participants (98%) with no prior preterm births and with measured birth weights and baseline hemoglobin values. Note 14 participants experienced stillbirth. Birth weights for these participants were included in the analysis; we return to this issue in the discussion. It was assumed that potential outcomes are independent of the exposure conditional on the following baseline covariates: trimester (1 or 2), maternal age group (18-24, 25-29, 30-34, 35-39, 40+), number of prior births (0, 1, 2, 3, 4+), alcohol use during pregnancy (yes or no), ART use (yes or no), and maternal height (cm) (Table 1). At baseline, no IPOP participants reported tobacco or drug use during pregnancy, so these covariates were not included. Primary analyses were conducted using R and were independently replicated in Python.
As discussed in the introduction, application of doubly robust estimators for estimating the $ACE$ of maternal anemia on birth weight may offer precision advantages over IPTW estimators, but the correct model specifications are unknown. To demonstrate differences in the influence function based variance estimator and the empirical sandwich variance estimator presented in Section 2.3, each of the doubly robust estimators were applied with three sets of model specifications. First, the covariates in Table 1 were included in both propensity and outcome models. Then, a naive outcome model was fit that included only the exposure variable (anemia status), while the propensity model included the full set of covariates. Under this specification, the outcome model was reasonably thought to be misspecified due to its simplicity. Finally, a naive propensity model was fit that included only an intercept term, while the outcome model included the full set of covariates. Due to the simplicity of the propensity model, it was reasonably thought to be misspecified. For all estimators, maternal height was modeled using restricted cubic splines with four knots placed at the 5th, 35th, 65th, and 95th percentiles (Perperoglou et al.,, 2019). The empirical sandwich variance estimator was computed using geex in R (Saul and Hudgens,, 2020) and delicatessen in Python (Zivich et al., 022c, ). Confidence intervals for the $ACE$ were constructed based on the influence function based variance estimator and the empirical sandwich variance estimator. Estimates for each of the three methods and model specifications along with corresponding 95% CIs are presented in Table 2, with 95% CI half-widths compared in Figure 1.
IPOP participants with and without anemia at baseline had similar self-reported alcohol use during pregnancy and similar age and height distributions. Participants without anemia were more likely to have one or more prior births, be in their first trimester, and be on ART. The three doubly robust methods provided similar estimates of the $ACE$ for all model specifications examined, with average birth weights if all women had anemia were approximately 40g lower than birth weights if no women had anemia, though all 95% CIs included zero.
Precision estimates based on the empirical sandwich variance estimator and the influence function based variance estimator were similar when the full covariate set was included in both the outcome and propensity models. Under the naive outcome model, standard errors for the influence function based variance estimator were larger than those for the empirical sandwich variance estimator, resulting in wider CIs. When the propensity model was naive, standard errors for the influence function based variance estimator were smaller than those for the empirical sandwich variance estimator, resulting in narrower CIs. Note there was more fluctuation in 95% CI half-widths across model specifications for the influence function based variance estimator than the empirical sandwich variance estimator (Figure 1).
Table 1: Baseline characteristics of IPOP participants by anemia status
No anemia $n=664$ Anemia $n=118$
Age category 18-24 156 (23%) 28 (24%)
25-29 190 (29%) 33 (28%)
30-34 192 (29%) 31 (26%)
35-39 100 (15%) 21 (18%)
40+ 26 (4%) 5 (4%)
Number of prior births 0 125 (19%) 30 (25%)
1 160 (24%) 25 (21%)
2 157 (24%) 28 (24%)
3 125 (19%) 18 (15%)
4+ 97 (15%) 17 (14%)
First trimester 120 (18%) 12 (10%)
ART use 648 (98%) 109 (92%)
Alcohol use during pregnancy 64 (10%) 9 (8%)
Height Median (Q1,Q3) 155 (150, 160) 154 (150, 159)
Mean (SD) 155 (8) 153 (8)
Min, Max 131, 185 132, 170
  • Note: Anemia defined as baseline hemoglobin below 10.5 g/dL
ACE ES-SE ES-95% CI IF-SE IF-95% CI
Full Covariate Set
Classic AIPW -37 56 (-147, 73) 58 (-151, 76)
Weighted regression AIPW -41 56 (-151, 69) 56 (-151, 68)
TMLE -37 56 (-147, 73) 58 (-151, 76)
Naive Outcome Model
Classic AIPW -36 57 (-148, 76) 61 (-156, 84)
Weighted regression AIPW -36 57 (-148, 77) 61 (-156, 84)
TMLE -36 57 (-148, 77) 61 (-156, 84)
Naive Propensity Model
Classic AIPW -41 58 (-154, 73) 56 (-150, 69)
Weighted regression AIPW -47 57 (-159, 65) 54 (-153, 59)
TMLE -40 58 (-154, 73) 56 (-150, 69)
  • Note: SE=standard error, ES=empirical sandwich, IF=influence function, CI=confidence interval. The naive outcome model included only the exposure, and the naive propensity model included only an intercept.
4 Simulation Study
A simulation study was conducted to compare the empirical properties of the influence function based variance estimator and the empirical sandwich variance estimator in conjunction with each of the three doubly robust methods discussed in Section 2.
4.1 Simulation Setup
Simulations were designed based on the example in Section 3, which aimed to estimate the effect of maternal anemia on birth weight among women with HIV. The primary simulations were conducted with $n=800$ participants, similar to the sample size in the example. Results for a sample size of $n=2000$ are included in the Supporting Information. The simulation study was conducted in R with code provided on GitHub.
Three covariates ($Z_{1}$, $Z_{2}$, $Z_{3}$), the exposure ($X$), and potential outcomes ($Y^{0}$ and $Y^{1}$) were simulated. The covariate $Z_{1}$ was distributed Normal with mean $155$ and standard deviation $7.6$. Two binary covariates $Z_{2}$ and $Z_{3}$ were simulated from Bernoulli distributions with means $0.25$ and $0.75$, respectively. The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(15-0.1Z_{1}+2.5Z_{2}-1Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$. Potential outcomes $Y^{0}$ and $Y^{1}$ under exposure and no exposure, respectively, were simulated from a normal distribution with mean $E(Y^{x})=1000+11.5Z_{1}+100Z_{2}-15Z_{1}Z_{2}+25x-5.5xZ_{1}-30xZ_{2}+20xZ_{1}Z_{2}$ and standard deviation $\sigma=400$ for $x\in\{0,1\}$. Under this data generating mechanism, the $ACE$ was approximately $-60$. Additional simulations were conducted with $\sigma=200$ with results in the Supporting Information.
The estimators $\widehat{DR}_{C}$, $\widehat{DR}_{WR}$, and $\widehat{DR}_{TMLE}$ were applied to 5000 simulated samples to estimate the $ACE$. Estimators were computed four ways: (1) with both propensity and outcome models correctly specified, (2) with propensity models correctly specified but outcome models misspecied, (3) with outcome models correctly specified but propensity models misspecified, and (4) with both models misspecified. Misspecified propensity models included only an intercept and a linear term for $(Z_{1}-155)^{2}$, and misspecified outcome models included only an intercept and linear terms for $X$ and $(Z_{1}-155)^{2}$.
[…]
4.2 Simulation Results
Both point and variance estimators performed as expected in simulations. The results of the simulation study for $n=800$ and $\sigma=400$ are presented in Table 3 and Figure 2. When at least one model was correctly specified, the classic AIPW, weighted regression AIPW, and TMLE displayed minimal bias, but all were substantially biased when both nuisance models were misspecified. Note bias was further reduced under correct specification of at least one model for $n=2000$ (Table 4). Under correct specification of both nuisance models, the empirical sandwich and influence function based variance estimators tracked closely with the ESE, resulting in SERs close to one and CIs attaining the nominal level of coverage. When both models were misspecified, both variance estimators tracked with the ESE, but bias was substantial, resulting in CIs with below nominal coverage.
Differences between the variance estimators are apparent when examining the simulation study results for scenarios where one nuisance model was correctly specified and the other was misspecified. When either the outcome model or the propensity model was misspecified, CIs based on the empirical sandwich variance estimator attained the nominal level of coverage, and estimated standard errors tracked closely with the ESE. In Figure 2, note variance ratios for the empirical sandwich estimator are closely clustered around one for these scenarios. In contrast, the influence function based variance estimator was empirically biased when either nuisance model was misspecified. As expected, it tended to overestimate the variance when the outcome model was misspecified, leading to estimated standard errors that tracked above the ESE, SERs above one, and conservative CIs that exceeded the nominal level of coverage. When the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator underestimated the variance, resulting in estimated standard errors below the ESE and SERs below one. The biases observed for the influence function based variance estimator were also present when $n=2000$ (Table 4 and Figure 3).
In the $\sigma=200$ scenarios, the empirical sandwich variance estimator performance was similar to the $\sigma=400$ scenario, but the performance of the influence function based variance estimator varied (Tables 5-6 and Figures 4-5). Specifically, when the propensity model was misspecified but the outcome model was correctly specified, the influence function based variance estimator demonstrated minimal empirical bias, resulting in SERs close to one.
In summary, the simulation study demonstrated the theoretical properties explained in Section 2. That is, the empirical sandwich variance estimator was consistent when at least one of the two nuisance models was correctly specified, and performed well in simulations. The influence function based variance estimator is not consistent under misspecification of either nuisance model. In line with expectations, the influence function based variance estimator was conservative when the outcome model was misspecified. For this data generating mechanism, it was anti-conservative when the propensity model was misspecified for $\sigma=400$ and exhibited minimal bias when $\sigma=200$.
Figure 2: Ratio between each simulation’s estimated standard error and the empirical standard error by estimator and model specification, n=800, $\sigma=400$, 5000 simulations. Black squares denote the mean variance ratio (=SER). Results exclude one simulation where models failed to converge. Note the $0.3\%$ of correct model specification simulations and $4.5\%$ of misspecified outcome model simulations where the ratio was above 1.2 are not displayed.
Table 3: Simulation summary results, $n=800$, $\sigma=400$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic 0.4 0.6 58.4 58.0 0.99 95 58.1 1.00 95
WR 0.4 0.6 58.3 57.7 0.99 95 57.6 0.99 95
TMLE 0.4 0.6 58.3 57.9 0.99 95 58.1 1.00 95
MO Classic -0.4 0.6 60.0 59.2 0.99 95 64.5 1.07 97
WR -1.7 2.8 59.2 58.4 0.99 95 62.4 1.05 96
TMLE -0.5 0.8 59.7 58.9 0.99 95 63.0 1.06 96
MW Classic 0.3 0.5 57.8 57.6 1.00 95 55.8 0.97 94
WR 0.2 0.4 57.7 57.6 1.00 95 55.8 0.97 94
TMLE 0.3 0.5 57.8 57.6 1.00 95 55.8 0.97 94
MB Classic -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
WR -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
TMLE -23.8 39.8 57.0 56.8 1.00 92 56.8 1.00 92
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%. Results exclude one simulation where models failed to converge.
5 Discussion
Doubly robust estimators have increased in popularity in the field of causal inference due to their ability to provide consistent point estimates when either an outcome or propensity model is correctly specified (Smith et al.,, 2023). In this paper, the commonly-used influence function based variance estimator is compared with the empirical sandwich variance estimator in conjunction with three doubly robust estimators: the classic AIPW estimator, the weighted regression AIPW estimator, and TMLE. These three estimators can be specified as solutions to sets of unbiased estimating equations, and thus their variances can be consistently estimated by the empirical sandwich variance estimator. Because the estimating equations remain unbiased when either the outcome or propensity model is correctly specified, the empirical sandwich variance estimator is a doubly robust variance estimator. In contrast, the influence function based variance estimator is consistent only when both outcome and propensity models are correctly specified. This discrepancy may be of particular importance in settings where the functional form of one of the nuisance models is known but the other is not. For example, with randomized trials, randomization to treatment means that the exposure mechanism is known. Doubly robust estimators can offer improved efficiency in this setting, but some of that efficiency gain may not be realized with application of the influence function based variance estimator, which is conservative if the outcome model is misspecified.
While consideration in this paper was limited to finite-dimensional parametric modeling approaches, machine learning approaches can also be used with doubly robust estimators (Kennedy et al.,, 2016; van der Laan and Gruber,, 2011, 2012; van der Laan et al.,, 2018; Balzer et al.,, 2023). Machine learning approaches allow for more complex functional forms for continuous covariates than fully parametric approaches and include higher order interaction terms that may not be included in investigator-specified parametric models, often leading to estimators that are more robust to model misspecification (Zivich et al.,, 2022). However, convergence of machine learning algorithms is typically slower and consistent variance estimation is more challenging than with parametric modeling approaches. Machine learning methods are typically not compatible with the estimating equations approach discussed in this paper, though alternative methods have been developed for doubly robust variance estimation in this context for TMLE estimators (Benkeser et al.,, 2017).
In the example, the effect of anemia on birth weight among women with HIV was estimated by applying the three doubly robust methods to data from the IPOP randomized trial. Average birth weights if all women had anemia were estimated to be approximately 40g lower than birth weights if no women had anemia, though all 95% CIs included zero. The analysis included 14 participants who experienced stillbirth. Because stillbirth is a competing outcome for live birth weight, future work could reexamine this problem using alternative approaches for competing events.
Use of the empirical sandwich variance estimator in conjunction with doubly robust estimators is not new, though its doubly robust property has not been emphasized. Lunceford and Davidian, (2004) discuss use of the empirical sandwich variance estimator for estimating the variance of the classic AIPW estimator and note that it tends to be more stable than the influence function based variance estimator. However, based on the systematic review discussed in the introduction (Smith et al.,, 2023), the influence function based variance estimator is frequently applied and the empirical sandwich variance estimator does not appear to be widely used with TMLE. While some existing software packages compute empirical sandwich variance estimators for doubly robust estimators (e.g., the “causaltrt” procedure in SAS and the “dr” procedure in Stata), other popular software packages compute the influence function based variance estimator but not the empirical sandwich variance estimator (e.g., the AIPW and tmle packages in R and the zEpid package in Python). We provide example R and Python code for estimating the variance of each of the three doubly robust estimators with the empirical sandwich variance estimator, which we hope will allow for easier implementation of doubly robust point and variance estimation.
Supplementary Material
Acknowledgements
This work was supported by US National Institutes of Health under award numbers R01 AI157758, R01 AI085073, P30 AI050410, and K01 AI177102. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. The authors thank Bradley Saul at the University of North Carolina at Chapel Hill for coding support.
Appendix
Consistency proof of TMLE estimator under correct specification of either the outcome or propensity model
@@ Acknowledgements
where $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$ is the vector of parameters.
Let
@@
Table 4: Simulation summary results, $n=2000$, $\sigma=400$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic -0.1 0.2 36.6 36.7 1.00 95 36.8 1.00 95
WR -0.1 0.2 36.6 36.6 1.00 95 36.6 1.00 95
TMLE -0.1 0.2 36.6 36.7 1.00 95 36.8 1.00 95
MO Classic -0.3 0.6 37.2 37.4 1.01 95 40.6 1.09 97
WR -0.9 1.6 37.0 37.1 1.00 95 39.6 1.07 96
TMLE -0.4 0.6 37.1 37.3 1.01 95 39.8 1.07 96
MW Classic -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 94
WR -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 95
TMLE -0.2 0.3 36.2 36.4 1.01 95 35.3 0.98 94
MB Classic -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
WR -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
TMLE -23.8 39.8 35.9 35.9 1.00 90 35.9 1.00 90
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%.
Table 5: Simulation summary results, $n=800$, $\sigma=200$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic 0.3 0.6 50.2 50.0 1.00 95 50.0 1.00 95
WR 0.3 0.5 50.1 49.9 1.00 95 49.9 1.00 95
TMLE 0.3 0.5 50.2 50.0 1.00 95 50.0 1.00 95
MO Classic -0.4 0.6 52.0 51.5 0.99 95 57.2 1.10 97
WR -1.6 2.7 51.3 50.7 0.99 95 55.1 1.07 96
TMLE -0.5 0.8 51.6 51.1 0.99 95 55.6 1.08 96
MW Classic 0.3 0.5 50.0 49.9 1.00 95 49.4 0.99 95
WR 0.3 0.4 50.0 49.9 1.00 95 49.4 0.99 95
TMLE 0.3 0.5 50.0 49.9 1.00 95 49.4 0.99 95
MB Classic -23.6 39.4 50.7 50.4 0.99 92 50.4 0.99 92
WR -23.6 39.4 50.7 50.4 0.99 92 50.4 0.99 92
TMLE -23.5 39.3 50.7 50.4 0.99 92 50.4 0.99 92
@@
  • ASE=average estimated standard error; ESE=empirical standard error; SER=standard error ratio (ASE/ESE); Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MW=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%. Results exclude 14 simulations where models did not converge.
@@
Table 6: Simulation summary results, $n=2000$, $\sigma=200$, $5000$ simulations. Bias, Relative Bias, ASE, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias Relative Bias (%) ESE ASE, ES SER, ES Cov, ES (%) ASE, IF SER, IF Cov, IF (%)
CS Classic 0.0 0.0 31.6 31.6 1.00 95 31.7 1.00 95
WR 0.0 0.1 31.5 31.6 1.00 95 31.6 1.00 95
TMLE 0.0 0.0 31.6 31.6 1.00 95 31.7 1.00 95
MO Classic -0.2 0.4 32.3 32.4 1.00 95 36.0 1.11 97
WR -0.8 1.4 32.1 32.1 1.00 95 34.9 1.09 97
TMLE -0.3 0.4 32.1 32.3 1.00 95 35.1 1.09 97
MW Classic 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
WR 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
TMLE 0.0 0.0 31.4 31.6 1.00 95 31.2 0.99 95
MB Classic -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
WR -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89
TMLE -23.6 39.5 32.0 31.9 1.00 89 31.9 1.00 89

### v2

@@ Acknowledgements
where $\psi_{\gamma}$ is the vector of score functions for the scaled outcome model. The estimating functions $\psi_{t1}$ and $\psi_{t0}$ are for the two targeting models, while $\psi_{\mu_{TMLE1}}$ and $\psi_{\mu_{TMLE0}}$ are for the causal means $\mu_{TMLE}^{1}$ and $\mu_{TMLE}^{0}$, respectively. Note $c_{x}(O_{i},\gamma,\eta_{x})=\mbox{expit}[\mbox{logit}\{a_{x}(Z_{i},\gamma)\}+\eta_{x}](b-a)+a$ are the pseudo-outcomes following the targeting step, as outlined in Section 2.2.3 of the main text. Finally, $\psi_{TMLE}$ is a delta method transformation of the causal means for estimation of the $ACE$ using $\widehat{DR}_{TMLE}$. Unbiasedness of $\psi_{3}(O_{i};\theta)$ when at least one working model is correctly specified is demonstrated in Appendix B.
Appendix B: Consistency proof of TMLE estimator under correct specification of either the outcome or propensity model
@@
where $\theta=[\alpha,\gamma,\eta_{1},\eta_{0},\mu_{TMLE}^{1},\mu_{TMLE}^{0},DR_{TMLE}]$ is the vector of parameters. Let
@@
Appendix C: Additional Simulation Scenarios
Additional Simulation Setup
In addition to simulations described in Section 4 of the main text, 10 additional simulation scenarios were considered:
  1. 1.
    continuous outcome, $n=800$, $\sigma=200$, $ACE\approx-60$
  2. 2.
    continuous outcome, $n=800$, $\sigma=600$, $ACE\approx-60$
  3. 3.
    continuous outcome, $n=2000$, $\sigma=200$, $ACE\approx-60$
  4. 4.
    continuous outcome, $n=2000$, $\sigma=400$, $ACE\approx-60$
  5. 5.
    continuous outcome, $n=2000$, $\sigma=600$, $ACE\approx-60$
  6. 6.
    continuous outcome, $n=2000$, $\sigma=400$, $ACE=0$ (null)
  7. 7.
    binary outcome, $n=800$, $ACE\approx 0.045$
  8. 8.
    binary outcome, $n=800$, $ACE=0$ (null)
  9. 9.
    binary outcome, $n=2000$, $ACE\approx 0.045$
  10. 10.
    binary outcome, $n=2000$, $ACE=0$ (null)
The continuous outcome scenarios followed the same data generating mechanism described in Section 4 of the main text, with the specified sample size and value of $\sigma$. For binary outcome scenarios, the distributions of $Z_{1}$, $Z_{2}$, and $Z_{3}$ were the same as the continuous outcome scenarios, but the exposure and outcomes differed. The exposure $X$ was simulated from a Bernoulli distribution with mean $\mbox{expit}(16-0.1Z_{1}+2.5Z_{2}-Z_{3}-0.02Z_{1}Z_{2}+0.005Z_{1}Z_{3})$. Potential outcomes $Y^{0}$ and $Y^{1}$ under no exposure and exposure, respectively, were simulated from Bernoulli distributions with means $\mbox{expit}\{-2+0.5(Z_{1}-155)+0.75Z_{2}-0.8Z_{2}(Z_{1}-155)-0.2(Z_{1}-155)(1-x)-0.1Z_{2}(1-x)+Z_{2}(Z_{1}-155)(1-x)\}$ for $x\in\{0,1\}$. Under this data generating mechanism, $ACE\approx 0.045$. To examine performance of the estimators under the null, simulations were also conducted for the binary outcome where $E(Y^{1})$ was equal to $E(Y^{0})$, as defined above, such that $ACE=0$.
Additional Simulation Results
Results of the additional simulation scenarios are presented in Appendix Tables A2-A11 and Appendix Figures A2-A11. When $n=2000$, empirical bias in point estimates, empirical sandwich variance estimates, and nonparametric bootstrap estimates under correct specification of at least one model was further reduced. However, an increase in the sample size did not mitigate bias in the influence function based variance estimator when one model was misspecified. For continuous outcomes, under misspecification of the outcome model, the conservative property of the influence function based variance estimator decreased as $\sigma$ increased. Under misspecification of the propensity model, influence function based variance estimators became increasingly anti-conservative as $\sigma$ increased.
For the binary outcome, the conservative property of the influence function based variance estimator under outcome model misspecification was more pronounced than for the continuous outcome. Notably, under misspecified propensity models, the influence function based variance estimator was slightly conservative when there was non-null effect but slightly anti-conservative under the null. These findings highlight that the direction and magnitude of bias under propensity model misspecification vary across data generating mechanisms.
Table A1: Baseline characteristics of IPOP participants by anemia status
No anemia $n=664$ Anemia $n=118$
Age category 18-24 156 (23%) 28 (24%)
25-29 190 (29%) 33 (28%)
30-34 192 (29%) 31 (26%)
35-39 100 (15%) 21 (18%)
40+ 26 (4%) 5 (4%)
Number of prior births 0 125 (19%) 30 (25%)
1 160 (24%) 25 (21%)
2 157 (24%) 28 (24%)
3 125 (19%) 18 (15%)
4+ 97 (15%) 17 (14%)
First trimester 120 (18%) 12 (10%)
ART use 648 (98%) 109 (92%)
Alcohol use during pregnancy 64 (10%) 9 (8%)
Height (cm) Median (Q1,Q3) 155 (150, 160) 154 (150, 159)
Mean (SD) 155 (8) 153 (8)
Min, Max 131, 185 132, 170
  • Note: Anemia defined as baseline hemoglobin below 10.5 g/dL
@@
Table A2: Simulation summary results, continuous outcome, $n=800$, $\sigma=200$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
CS Classic 0.4 50.1 1.00 95 1.00 95 1.00 95
WR 0.4 50.1 1.00 95 1.00 95 1.00 95
TMLE 0.3 50.1 1.00 95 1.00 95 1.00 95
MO Classic -0.3 51.9 0.99 95 1.02 96 1.10 97
WR -1.6 51.3 0.99 95 1.00 95 1.07 96
TMLE -0.4 51.6 0.99 95 1.01 95 1.08 96
MP Classic 0.4 50.0 1.00 95 1.00 95 0.99 95
WR 0.3 50.0 1.00 95 1.00 95 0.99 95
TMLE 0.3 50.0 1.00 95 1.00 95 0.99 95
MB Classic -23.5 50.7 0.99 92 1.00 92 0.99 92
WR -23.5 50.7 0.99 92 1.00 92 0.99 92
TMLE -23.5 50.7 0.99 92 1.00 92 0.99 92
@@
  • ESE=empirical standard error; SER=standard error ratio (ASE/ESE), where ASE=average estimated standard error; Cov = 95% confidence interval coverage; ES=empirical sandwich variance estimator; NB=nonparametric bootstrap variance estimator, IF=influence function based variance estimator; CS=correct specification of both models; MO=misspecified outcome model, MP=misspecified propensity model; MB=misspecified both models; WR=weighted regression AIPW; $ACE$ was approximately $-60$; Monte Carlo standard error for 95% CI coverage was 0.3% when coverage was 95%. Results exclude 15 simulations where models failed to converge.
@@
Table A3: Simulation summary results, continuous outcome, $n=800$, $\sigma=600$, $5000$ simulations. Bias, ESE, SER, and 95% CI coverage calculated for the ACE.
Scenario Estimator Bias ESE SER, ES Cov, ES (%) SER, NB Cov, NB (%) SER, IF Cov, IF (%)
CS Classic 0.4 70.1 0.99 94 1.00 95 0.99 95
WR 0.4 69.9 0.98 94 1.00 95 0.98 94
TMLE 0.4 70.1 0.99 94 1.00 95 0.99 95
MO Classic -0.4 71.5 0.98 95 1.01 95 1.05 96
WR -1.7 70.6 0.98 95 1.00 95 1.03 96
TMLE -0.5 71.2 0.98 95 1.00 95 1.04 95
MP Classic 0.3 68.9 0.99 95 1.00 95 0.95 93
WR 0.2 68.9 0.99 95 1.00 95 0.95 93
TMLE 0.3 68.9 0.99 95 1.00 95 0.95 93
MB Classic -24.1 66.2 1.00 93 1.00 93 1.00 93
WR -24.1 66.2 1.00 93 1.00 93 1.00 93
TMLE -24.1 66.2 1.00 93 1.00 93 1.00 93